Solve the radical expression
Please help with this!
Sqrt(2x-1) = x - 2
What is in parenthesis is under the sqrt symbol.

Answer by checkley71(8403)

sqrt(2x-1)=x-2 or squaring both sides we get 2x-1=(x-2)~2 or 2x-1=x~2-4x+4 or
x~2-6x+5=0 or (x-5)(x-1)=0 or x=5 & x=1